# Good News for Polynomial Root-finding

### v35

@@ 3.3 Approximation of the power sums in the unit disc
Cor. 3.1 implies that the values $|s_{h,q-}s_{h}|$ decrease exponentially in $q-h$ if the unit disc $D(0,1)$ is isolated.
@@
[130, Eqn. (12.10)]. [^20] Unlike the proof in [130], we rely on Cor. 3.1 of independent interest. Let the unit disc $D(0,1)$ be $\theta$-isolated and let $d_{\rm in}$ and $d_{\rm out}=d-d_{\rm in}$ denote the numbers of the roots lying in and outside the disc $D(0,1)$, respectively. Then
@@
Enumerate the roots $z_{j}$ so that $|z_{j}|<1$ if and only if $j\leq d_{\rm in}$; then Cor. 3.1 implies that
@@
Let the unit circle $C(0,1)$ be 2-isolated. Then Thm. 3.1 implies that $|s_{0,q}-s_{0}|<1/2$ for $d\leq 1,000$ if $q=10$ and for $d\leq 1,000,000$ if $q=20$.
@@
Observation 3.1.
(i) Alg. 3.1 evaluates $p(x)$ at $q=\lfloor\log_{\theta}(4d+2)\rfloor$ points. (ii) It outputs $\bar{s}_{0}=\#(D(0,1))$ if the circle $C(0,1)$ is $\theta$-isolated for $\theta>1$. (iii) If the algorithm outputs $\bar{s}_{0}>0$, then #$\#(D(0,\theta))>0$.
@@
Thm. 3.1 immediately implies claim (ii) but also implies that #$\#(D(0,1))>0$ if $\bar{s}_{0}>0$ unless #$\#(A(0,1/\theta,\theta))>0$. In both cases #$\#(D(0,\theta))>0$. ∎
@@ 3.4 Extension to any disc
Map (1.4), preserving #$\#(D(0,1))$ (cf. Observation 2.1), enables us to extend definition (3.2) of Cauchy sum to any disc $D(c,\rho)$ as follows:
@@
[…]
@@
[…]
By applying Thm. 3.1 we can estimate the errors $|s_{h,q}(t,0,1)-s_{h}(t,0,1)|$ of the approximation of the $h$th power sums $s_{h}(t,0,1)$ of the roots of $t(x)$ in the unit disc $D(0,1)$. Next we specify this for $h=1$.
@@ 3.5 The real part of the 0-th Cauchy sum
Recall that $s_{0,q}=\sum_{j=1}^{d}\frac{1}{1-z_{j}^{q}}$ unless $z_{j}^{q}=1$ for some $q$ (cf. Cor. 3.1), write $z:=z_{j}^{q}$ and estimate the contribution of a roots $z_{j}$ to $\Re(s_{0,q})$.
@@
Under the assumptions of Prop. 3.3 it holds that
@@
Next we outline but do not elaborate upon some potential links of Cor. 3.2 to root-counting. Let the unit disc $D(0,1)$ be $\theta$-isolated, combine bounds (3.15) and (3.16), and obtain
@@
Combine Cor. 3.1 for $h=0$ with equation (3.12). ∎
@@ 4.1 Overview
We seek fast root-counters and $\ell$-tests based on computing Cauchy sum $s_{0,q}$ for a smaller integer $q$, but the output can deceive us if we decrease $q$ too much: the polynomial $p(x):=(1-wx^{d-m})(x^{m}-dx+w)$ has zero near the origin if $w\approx 0$, while NIR$(x)=p^{\prime}(x)/p(x)$ vanishes for $x$ equal to any $(d-1)$st root of unity.
Such deceptive e/i tests have been extremely rare guests in the extensive experiments in [57, 60] for $q$ of order $\log^{2}(d))$ and were completely excluded where the values $|s_{h,q}|$ were required to be small for $h=0,1,2$, but formal support for this empirical behavior is still a challenge.
@@
One can probably obtain similar results by computing the Cauchy sum $s_{0,q}$ for the polynomial $p(ax)$ at $O(m\log(d))$ random values $a\in C(0,1)$, but we leave formal support of this conjecture as a research challenge.
Recall that randomization in our root-finders is of Las Vegas type (see Sec. 1.7), while we apply randomization in Sec. 6 anyway – to bound computational precision in our root-finders. For the sake of completeness we also present a deterministic $\ell$-test in Sec. 4.4.
@@ 4.2 A basic randomized root-counter
Alg. 4.1 queries evaluation oracle at $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$ locations and outputs the value $\bar{s}_{0}=\#(D(0,\rho))$ with a probability at least $1-1/\gamma$.
@@
Hence with a probability at least $1-1/\gamma$ the circle $C(0,\rho)$ is $\theta$-isolated for $\theta=2^{\frac{1}{10m\gamma}}$, in which case Alg. 4.1 outputs $\bar{s}_{0}=s_{0}=\#(D(0,\rho))$ by virtue of claim (ii) of Prop. 3.1. ∎
4.3 Refining the bound on error probability
By reapplying Alg. 4.1 $v$ times we evaluate the polynomial $p$ at $v$ times more points, but we will prove that this decreases the Las Vegas error probability dramatically – below $1/\gamma^{v}$.
First let $\ell=1$ or $\ell=m$ and then again reduce the root radius approximation problem to the decision problem of $\ell$-test. Namely, narrow Alg. 4.1 to an $\ell$-test for a fixed $\ell$ in the range $1\leq\ell\leq m$, so that with a probability at least $1-1/\gamma$ this $\ell$-test outputs 0 if #$\#(D(0,1))<\ell$ and outputs 1 if #$\#(D(0,2^{0.5}))\geq\ell$. Moreover, output 1 is certified by virtue of claim (iii) of Prop. 3.1 if $\ell=1$; similarly the output 0 is certified if $\ell=m$.
Next extend the $\ell$-test by means of applying Alg. 4.1 for $v$ iid random variables $w$ in the range $0.2\leq w\leq 0.4$. Call this $\ell$-test Alg. $\ell$.
@@ 4.3 Refining the bound on error probability
Recall Prop. 4.1 and then readily verify the following proposition.
@@
(i) For two integers $1\leq\ell\leq m$ and $v\geq 1$ and a real $\gamma\geq 1$, Alg. $\ell$ evaluates $p(x)$ at $\lfloor 10m\gamma\log_{2}(4d+2)\rfloor v$ points; this evaluation involves more ops than the other stages of the algorithm, even in the cases of the Mandelbrot and sparse polynomials $p$ (cf. Observation 3.2). (ii) Output 1 of Alg. 4.1v,1 and the output 0 of Alg. 4.1v,$m$ are certified. (iii) Output 0 of Alg. 4.1v,1 and output 1 of Alg. 4.1v,$m$ are correct with a probability at least $1-1/\gamma^{v}$.
@@
This extension of Alg. $\ell$ is only interesting for $1<\ell<m$; otherwise its cost/error estimates are inferior to those of Prop. 4.2.
@@
4.4 Deterministic $\ell$-test
In view of our remark at the end of Sec. 4.1, randomization is only a nominal restriction for $\ell$-tests, but next we get rid of it, although at the price of increasing the numbers of points of evaluation of $p$ and bit operations involved in the tests by a factor of $m$.
[…]
Next we elaborate upon this test.
Without loss of generality (wlog) let $c=0$, $\theta\rho=1$, $C(c,\theta\rho)=C(0,1)$.
Algorithm 4.2.
Deterministic $\ell$-test.
INPUT: A black box polynomial $p$ of a degree $d$ and two positive integers $\ell$ and $m\geq\ell$ such that the unit disc $D(0,1)$ contains at most $m$ roots.
INITIALIZATION: Fix
$\sigma>1,~{}\theta=\sigma^{\frac{1}{4m+1}},~{}q=\lfloor\log_{\theta}(4d+2)\rfloor,~{}\rho_{i}=\frac{1}{\theta^{2i+1}},~{}i=0,1,\dots,2m.$ (4.1)
OUTPUT: 0 if #$\#(D(0,1/\sigma))<\ell$ or 1 if #$\#(D(0,1))\geq\ell$.
[…]
Output 1 if at least $m+1$ integers $\bar{s}_{0}$ output by the algorithm (among all $12m+$ 1 such integers) exceed $\ell-1$. Otherwise output 0. (For $\ell=1$ we can output 1 already on a single output $\bar{s}_{0}>0$; for $\ell=m$ we can output 0 already on a single output $\bar{s}_{0}<m$.)
Proposition 4.4.
Alg. 4.2 is a $\sigma$-soft $\ell$-test for $\sigma>1$ of our choice (see (4.1)).
Proof.
[…]
To complete the proof, combine this property with the relationships $\rho_{2m}=1/\theta^{4m+1}$ (cf. (4.1)), $\rho_{0}=1/\theta<1$, and #$\#(D(0,\rho_{i+1}))\leq\#(D(0,\rho_{i}))$, which hold because $D(0,\rho_{i+1})\subseteq(D(0,\rho_{i})$ for $i=0,1,\dots,2m-1$. ∎
Observation 4.1.
[Cf. Observation 3.2.] Alg. 4.2 evaluates polynomial $p(x)$ at $(2m+1)q$ points for $q=\lfloor\log_{\theta}(4d+2)\rfloor=\lfloor(4m+1)\log_{\sigma}(4d+2)\rfloor$, which dominates the cost of the remaining ops involved, even in the cases of the Mandelbrot and sparse input polynomials and which is in $O(m^{2}\log(d))$ if $\sigma-1$ exceeds a positive constant.
@@ 5 Root radius estimation based on $\ell$-test
In this section we assume that we are given an algorithm for $\ell$-test and go back from the decision problem of $\ell$-test to approximation of $r_{d-\ell+1}(c^{\prime})$, the $\ell$th smallest root radius from a fixed complex point $c^{\prime}$; we assume that $r_{d-\ell+1}(c^{\prime})$ lies in a fixed range $[a_{-},a_{+}]$. [^22] For our root-finders we only need to approximate $r_{d-\ell+1}(c^{\prime})$ for $\ell=1$ and $\ell=m$ and only with a constant bound on the relative error, but we will estimate $r_{d-\ell+1}(c^{\prime})$ within $1/2^{b}$ for any real $b$ and $\ell$, $1\leq\ell\leq m$, which is of independent interest for root-finding. Wlog let this be the range $[\epsilon,1]$ for a fixed small positive $\epsilon=1/2^{b}$. The bisection algorithm is too slow for us, and we accelerate it by applying it to approximate the exponent of a target value. Call that algorithm bisection of exponent or BoE and apply it to bracket the $\ell$th smallest root radius $r_{d-\ell+1}$ in a range $[\rho_{-},\rho_{+}]$ such that $0<\rho_{<}-\rho_{+}\leq\epsilon$ or $0<\rho_{+}<\beta\rho_{-}$ for two fixed tolerance values $\epsilon=1/2^{b}$ and $\beta=2^{\phi}>1$.
[…]
Since the test is soft, so is BoE as well: for a range $[b_{-},b_{+}]$ with midpoint $\widehat{b}=0.5(b_{+}+b_{-})$ regular rigid bisection continues the search in the range $[b_{-},\widehat{b}]$ or $[\widehat{b},b_{+}]$, but our soft BoE continues it in the range $[b_{-},(1+\alpha)\widehat{b}]$ or $[(1-\alpha)\widehat{b},b_{+}]$ for a fixed small positive $\alpha$. The approximations still converges linearly and just a little slower than the regular rigid bisection: the approximation error decreases by a factor of $\frac{2}{1+\alpha}$ rather than by twice. [^23] If we applied BoE to the range $[0,1]$, we would have to bisect the infinite range $[-\infty,0]$, but we deal with the narrower range $[-b,0]$ since we have narrowed the range $[0,1]$ to $[\epsilon,1]$.
@@
[…]
@@
Readily verify that our policy of recursively updating the range $[b_{-},b_{+}]$ maintains this property throughout and that updating stops where $b_{+}-b_{-}\leq\phi$, that is, where $\rho_{+}\leq\beta\rho_{-}$. This implies correctness of Alg. 5.1.
@@
Alg. 5.1 invokes an $\ell$-test $\mathbb{I}$ times where $\mathbb{I}\leq\lceil\log_{\nu}(\frac{\bar{b}}{\phi})\rceil+1$ for $\nu=\frac{2}{1+\alpha}$, $\mathbb{I}=O(\log(b/\phi))$.
@@
The computations in Alg. 5.1 stop after the first application of an $\ell$-test if it outputs 1. It remains to count the subsequent $\ell$-tests if it outputs 0.
@@
Clearly, $\mathbb{I}$ applications of $\ell$-test increase the number of evaluation points for the ratio $\frac{p^{\prime}(x)}{p(x)}$ by at most a factor of $\mathbb{I}$, and if the $\ell$-test is randomized, then the error probability also grows by at most a factor of $\mathbb{I}$. Let us specify these estimates for Alg. 5.1.
@@
Let precisely $m$ zeros of a black box polynomial $p(x)$ of a degree $d$ lie in the unit disc, that is, let #$\#(D(0,1))=m$, and let $i(D(0,1))\geq\theta$, for $\theta-1$ exceeding a positive constant. Given two positive tolerance values $\epsilon=1/2^{b}$ and $\phi$, an integer $\ell$, $1\leq\ell\leq m$, and a black box $\ell$-test, which evaluates polynomial $p(x)$ at $\mathcal{NR}$ points, Alg. 5.1 approximates the $\ell$-th smallest root radius $|z_{j}|$ within the relative error $2^{\phi}$ or determines that $|z_{j}|\leq\epsilon$, at the cost of evaluation of polynomial $p(x)$ at $\mathbb{I}\cdot\mathcal{NR}$ points for $\mathbb{I}=O(\log(\frac{b}{\phi}))$ estimated in Prop. 5.1.

### v36

@@ 7.3 Approximation of the power sums in the unit disc
Cor. 7.1 implies that the values $|s_{h,q-}s_{h}|$ decrease exponentially in $q-h$ if the unit disc $D(0,1)$ is isolated.
@@
[134, Eqn. (12.10)]. [^12] Unlike the proof in [134], we rely on Cor. 7.1 of independent interest. Let the unit disc $D(0,1)$ be $\theta$-isolated and let $d_{\rm in}$ and $d_{\rm out}=d-d_{\rm in}$ denote the numbers of the roots lying in and outside the disc $D(0,1)$, respectively. Then
@@
Enumerate the roots $z_{j}$ so that $|z_{j}|<1$ if and only if $j\leq d_{\rm in}$; then Cor. 7.1 implies that
@@
Let the unit circle $C(0,1)$ be 2-isolated. Then Thm. 7.1 implies that $|s_{0,q}-s_{0}|<1/2$ for $d\leq 1,000$ if $q=10$ and for $d\leq 1,000,000$ if $q=20$.
@@
Observation 7.1.
(i) Alg. 7.1 evaluates $p(x)$ at $q=\lfloor\log_{\theta}(4d+2)\rfloor$ points. (ii) It outputs $\bar{s}_{0}=\#(D(0,1))$ if the circle $C(0,1)$ is $\theta$-isolated for $\theta>1$. (iii) If the algorithm outputs $\bar{s}_{0}>0$, then #$\#(D(0,\theta))>0$.
@@
Thm. 7.1 immediately implies claim (ii) but also implies that #$\#(D(0,1))>0$ if $\bar{s}_{0}>0$ unless #$\#(A(0,1/\theta,\theta))>0$. In both cases #$\#(D(0,\theta))>0$. ∎
@@ 7.4 Extension to any disc
Map (1.3), preserving #$\#(D(0,1))$ (cf. Observation 6.1), enables us to extend definition (7.2) of Cauchy sum to any disc $D(c,\rho)$ as follows:
@@
Refer to Alg. 7.1 applied to the polynomial $t(y)$ of (1.3) for an isolation parameter $\theta>1$ and $q=\lfloor\log_{\theta}(4d+2)\rfloor$ as Alg. 7.1c,ρ,q and also as Alg. 7.1 applied to $\theta$-isolated circle $C(c,\rho)$ for $2q$ evaluation points. Combine Observations 6.1 and 7.2 to extend Observation 7.1 as follows.
@@
(i) Alg. 7.1c,ρ,θ evaluates polynomial $p(x)$ at $q=\lfloor\log_{\theta}(4d+2)\rfloor$ points $x$ and in addition performs $25q-$ 2 ops. (ii) It outputs $\bar{s}_{0}=\#(D(c,\rho))$ if the circle $C(c,\rho)$ is $\theta$-isolated. (iii) If the algorithm outputs $\bar{s}_{0}>0$, then #$\#(D(c,\theta\rho))>0$.
By applying Thm. 7.1 we can estimate the errors $|s_{h,q}(t,0,1)-s_{h}(t,0,1)|$ of the approximation of the $h$th power sums $s_{h}(t,0,1)$ of the roots of $t(x)$ in the unit disc $D(0,1)$. Next we specify this for $h=1$.
@@ 7.5 The real part of the 0-th Cauchy sum
Recall that $s_{0,q}=\sum_{j=1}^{d}\frac{1}{1-z_{j}^{q}}$ unless $z_{j}^{q}=1$ for some $q$ (cf. Cor. 7.1), write $z:=z_{j}^{q}$ and estimate the contribution of a roots $z_{j}$ to $\Re(s_{0,q})$.
@@
Under the assumptions of Prop. 7.3 it holds that
@@
Next we outline but do not elaborate upon some potential links of Cor. 7.2 to root-counting. Let the unit disc $D(0,1)$ be $\theta$-isolated, combine bounds (7.15) and (7.16), and obtain
@@
Combine Cor. 7.1 for $h=0$ with equation (7.12). ∎
@@ 8.1 Overview
We seek fast root-counters and $\ell$-tests based on computing Cauchy sum $s_{0,q}$ for a smaller integer $q$, but the output can deceive us if we decrease $q$ too much: the polynomial $p(x):=x^{d}-dx+w$ has zero near the origin if $w\approx 0$, while NIR$(x)=p^{\prime}(x)/p(x)$ vanishes for $x$ equal to any $(d-1)$st root of unity.
Such deceptive e/i tests have been very rare guests in the extensive experiments in [60, 63] for $q$ of order $\log^{2}(d))$ and were completely excluded where the values $|s_{h,q}|$ were required to be small for $h=0,1,2$, but formal support for this empirical behavior is still a challenge.
@@
Combine this estimate with Cor. 2.2 for $k$ of order $b$ and obtain Las Vegas randomized solution of Problems 0, 1, 0, and 1 at the cost of evaluation of the polynomial $p$ at $O(m^{2}b)$ points.
@@ 8.2 A basic randomized root-counter
Alg. 8.1 queries evaluation oracle at $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$ locations and outputs the value $\bar{s}_{0}=\#(D(0,\rho))$ with a probability at least $1-1/\gamma$.
@@
Hence with a probability at least $1-1/\gamma$ the circle $C(0,\rho)$ is $\theta$-isolated for $\theta=2^{\frac{1}{10m\gamma}}$, in which case Alg. 8.1 outputs $\bar{s}_{0}=s_{0}=\#(D(0,\rho))$ by virtue of claim (ii) of Prop. 7.1. ∎
8.3 Refining the bound on the error probability
When we apply Alg. 8.1 $v$ times, the number of query locations of the evaluation oracle grows linearly – by a factor of $v$, while the bound on Las Vegas error probability of outputting errors decreases exponentially – from $1/\gamma$ to $1/\gamma^{v}$.
First let $\ell=1$ or $\ell=m$ and then again reduce the root radius approximation problem to the decision problem of $\ell$-test. Namely, narrow Alg. 8.1 to an $\ell$-test for a fixed $\ell$ in the range $1\leq\ell\leq m$, so that with a probability at least $1-1/\gamma$ this $\ell$-test outputs 0 if #$\#(D(0,1))<\ell$ and outputs 1 if #$\#(D(0,2^{0.5}))\geq\ell$. Moreover, output 1 is certified by virtue of claim (iii) of Prop. 7.1 if $\ell=1$; similarly the output 0 is certified if $\ell=m$.
Next extend the $\ell$-test by means of applying Alg. 8.1 for $v$ iid random variables $w$ in the range $0.2\leq w\leq 0.4$. Call this $\ell$-test Alg. $\ell$.
@@ 8.3 Refining the bound on the error probability
Recall Prop. 8.1 and then readily verify the following proposition.
@@
(i) For two integers $1\leq\ell\leq m$ and $v\geq 1$ and a real $\gamma\geq 1$, Alg. $\ell$ evaluates $p(x)$ at $\lfloor 10m\gamma\log_{2}(4d+2)\rfloor v$ points; this evaluation involves more ops than the other stages of the algorithm, even in the cases of the Mandelbrot and sparse polynomials $p$ (cf. Observation 7.2). (ii) Output 1 of Alg. 8.1v,1 and the output 0 of Alg. 8.1v,$m$ are certified. (iii) Output 0 of Alg. 8.1v,1 and output 1 of Alg. 8.1v,$m$ are correct with a probability at least $1-1/\gamma^{v}$.
@@
This extension of Alg. $\ell$ is only interesting for $1<\ell<m$; otherwise its cost/error estimates are inferior to those of Prop. 8.2.
@@ 9 Root radius estimation based on $\ell$-test
In this section we assume that we are given an algorithm for $\ell$-test and seek approximation of $r_{d-\ell+1}(c^{\prime})$, the $\ell$th smallest root radius from a fixed complex point $c^{\prime}$, known to lie in a fixed range $[a_{-},a_{+}]$. [^14] For our root-finders we only need to approximate $r_{d-\ell+1}(c^{\prime})$ for $\ell=1$ and $\ell=m$, but we estimate $r_{d-\ell+1}(c^{\prime})$ for any $\ell$, $1\leq\ell\leq m$, which is of independent interest for root-finding.
Wlog let this be the range $[\epsilon,1]$ for a fixed small positive $\epsilon=1/2^{b}$ and seek a range $[\rho_{-},\rho_{+}]$ such that $0\leq\rho_{-}\leq r_{d-\ell+1}(c^{\prime})\leq\rho_{+}$ and either $\rho_{+}\leq\epsilon$ or else $0<\rho_{+}<\beta\rho_{-}$ for two fixed tolerance values $\epsilon=1/2^{b}$ and $\beta=2^{\phi}>1$. We can obtain the desired approximation by applying bisection algorithm to the range $[\epsilon,1]$, but this is too slow for us, and we apply bisection of logarithms or BoL to approximate the logarithm of the target value in the range $[-b,0]=[\log_{2}(\epsilon),\log_{2}(1)]$.
Assume that besides the above tolerance values the input of BoL includes a softness bound $\sigma=1+\alpha$, for a fixed small positive parameter $\alpha<1/2$, and a $\sigma$-soft $\ell$-testc,ρ, to be used as the pivot of BoL.
Since the test is soft, so is BoL as well: for a range $[b_{-},b_{+}]$ with midpoint $\widehat{b}=0.5(b_{+}+b_{-})$ regular rigid bisection continues the search in the range $[b_{-},\widehat{b}]$ or $[\widehat{b},b_{+}]$, but our soft BoL continues it in the range $[b_{-},(1+\alpha)\widehat{b}]$ or $[(1-\alpha)\widehat{b},b_{+}]$ for a fixed small positive $\alpha$. Accordingly we slightly increase range of search for $r_{d-\ell+1}(c^{\prime})$ to $[-\bar{b},0]$ for $\bar{b}:=\log_{2}(\frac{\sigma}{\epsilon})$. BoL still converges linearly and just a little slower than the regular rigid bisection: the approximation error decreases by a factor of $\frac{2}{1+\alpha}$ rather than by twice. [^15] If we applied BoL to the range $[0,1]$, we would have to bisect the infinite range $[-\infty,0]$, but we deal with the narrower range $[-b,0]$ since we have narrowed the range $[0,1]$ to $[\epsilon,1]$.
@@
INPUT: Two integers $\ell$ and $m$ such that $1\leq\ell\leq m$ and #$\#(D(0,1))\leq m$, $\sigma=1+\alpha$ for $20<\alpha<1/$ 2, a $\sigma$-soft $\ell$-test0,ρ, and two tolerance values $\beta:=2^{\phi}>1$ and $\epsilon=1/2^{b}$, $b\geq 1$.
@@
Readily verify that our policy of recursively updating the range $[b_{-},b_{+}]$ maintains this property throughout and that updating stops where $b_{+}-b_{-}\leq\phi$, that is, where $\rho_{+}\leq\beta\rho_{-}$. This implies correctness of Alg. 9.1.
@@
Alg. 9.1 invokes an $\ell$-test $\mathbb{I}$ times where $\mathbb{I}\leq\lceil\log_{\nu}(\frac{\bar{b}}{\phi})\rceil+1$ for $\nu=\frac{2}{1+\alpha}$, $\mathbb{I}=O(\log(b/\phi))$.
@@
The computations in Alg. 9.1 stop after the first application of an $\ell$-test if it outputs 1. It remains to count the subsequent $\ell$-tests if it outputs 0.
@@
Clearly, $\mathbb{I}$ applications of $\ell$-test increase the number of evaluation points for the ratio $\frac{p^{\prime}(x)}{p(x)}$ by at most a factor of $\mathbb{I}$, and if the $\ell$-test is randomized, then the error probability also grows by at most a factor of $\mathbb{I}$. Let us specify these estimates for Alg. 9.1.
@@
Let precisely $m$ zeros of a black box polynomial $p(x)$ of a degree $d$ lie in the unit disc, that is, let #$\#(D(0,1))=m$, and let $i(D(0,1))\geq\theta$, for $\theta-1$ exceeding a positive constant. Given two positive tolerance values $\epsilon=1/2^{b}$ and $\phi$, an integer $\ell$, $1\leq\ell\leq m$, and a black box $\ell$-test, which evaluates polynomial $p(x)$ at $\mathcal{NR}$ points, Alg. 9.1 approximates the $\ell$-th smallest root radius $|z_{j}|$ within the relative error $2^{\phi}$ or determines that $|z_{j}|\leq\epsilon$, at the cost of evaluation of polynomial $p(x)$ at $\mathbb{I}\cdot\mathcal{NR}$ points for $\mathbb{I}=O(\log(\frac{b}{\phi}))$ estimated in Prop. 9.1.